Operations Report – March 20th

Crew 207 Operations Report 20 Mar 2019

SOL: 4

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator:Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 11:00 hours

Generator was turned off at 07:30, SOC 100%

Diesel Reading – 53 %

Propane Reading – 68 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 250 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 164 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 141939.6 gallons

Toilet tank emptied: Yes

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:53.6 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Curiosity rover used: Not used

Hours: 84.4 hours

Beginning charge: 100 %

Ending charge: 100%

Currently charging: Yes

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where?

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Power System Non-nominal, drop down to SOC 0%, after turning the generator, SOC went to 100%. Current SOC% 95%, generator turned on at 20:00

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Astronomy Report – March 20th

Astronomy Report

Name: Tom Baldwin

Crew: 207

Date: 20-MAR-2019

MDRS ROBOTIC OBSERVATORY

Robotic Telescope MDRS-WF: processed images of Horsehead Nebula, (attached)

Objects Viewed: none

Problems Encountered: none

MUSK OBSERVATORY : Viewed Sun through eyepiece and camera

Solar Features Observed: Sun

Problems Encountered: none

Captures to be processed

EVA Report – March 20th

Crew 207 EVA Report 20 Mar 2019

EVA #3

Author : Julielynn Wong

Purpose of EVA: Allow Part 107 novice drone pilots to (1) fly a Mavic Pro drone with a standard controller in a symmetric figure of 8 flight pattern; (2) prepare to conduct an outdoor inspection of the MDRS habitat roof under VLOS in a simulated emergency drill

Start time: 14:40

End time: 16:10

Narrative: Software issues led to the amendment of today’s EVA plan. We modified our plan to move ahead on drone flight testing that was planned for later on this mission.

The weather was clear and sunny. We walked halfway to North Ridge (4251500 N / 518250 W).

A demo flight of VTOL, pitch, yaw, roll, and autonomous landing was piloted by XO using the Mavic Pro in Mode 2 stick configuration. ENG then flew the drone, familiarized himself with the flight controls, and attempted to perform a symmetric figure-of-8 while maintaining altitude, and landed the drone. HSO then flew the drone, familiarized herself with the flight controls, and attempted to perform a symmetric figure-of-8 while maintaining altitude, and landed the drone. The drone remained stable despite high gusting winds present during the entire flight.

ENG and GHO then collected soil samples for analysis using soil sieves 3D printed by SCI and HSO. XO performed a test flight for a bird’s eye view inspection of the MDRS habitat for a planned emergency simulation.

We identified the following useful items to 3D print for an EVA at MDRS:

Headset ear piece

Destination: Halfway to North Ridge

Coordinates: Halfway to 44251500 N / 518250 W (North Ridge); our GPS device ran out of batteries and will be replaced by XO

EVA Participants: Dean (XO), Tom Baldwin (ENG), Diane Rothberg (GHO)

Roads and routes per MDRS Map: Road 1103

Mode of travel: Walking

Operations Report – March 20th

Crew 207 Operations Report 20 Mar 2019

SOL: 4

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator-Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 11:00 hours

Generator was turned off at 07:30, SOC 100%

Diesel Reading – 53 %

Propane Reading – 68 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 250 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 164 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 141939.6 gallons

Toilet tank emptied: Yes

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 66 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:53.6 hours

Beginning charge: 100 %

Ending charge: 56 %

Currently charging: Yes

Curiosity rover used: Not used

Hours: 84.4 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where?

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Greenhab Report – March 20th

Crew 207 GreenHab Report – 20 Mar 2019

GreenHab Officer: Diane Rothberg

Environmental control: Heat and cooling w/ambient air (8:35))

Shade Cloth (40%)

Average temperatures:

Low: 16.0 deg C

High: 25 deg C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 5.5 gallons plus 1 extra gal to hydrate soil for new crops of radishes and cucumbers. Total: 6.5 gal

Water in Blue Tank: 163.5 gallons

Crops watered at 09:55 hours (2 gal) and 17:30 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 9:55 AM 22 deg C / 20%

2) 13:00 AM 23 deg C / 20%

3) 16:35 PM 21 deg C / 20%

4) 18:15 PM 17 deg C / 20%

Changes to crops: New sprouts on Kale and spinach planted 3 days ago

Narrative: Used an extra gallon of water to hydrate dry soil for planting new tray of radishes and new pot of cucumbers. It may look like a lot, but we are only harvesting where crop is abundant.

Harvest: Basil: 23 gm, cilantro .05 gm, oregano .05 gm

Support/ Supplies needed: none

Sol Summary – March 20th

Summary Title : Day #4 of Sim

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary: GHO test rode the bike-powered 3D printing apparatus for 1 hour in Level 1 of the Hab last night. Last night and today, COM collected study data for the thoracentesis project in the Science Dome. Crew 207 continued 3D printing systems in Level 1 of the Hab. AST continued operations in the solar observatory. Weather was clear and sunny with gusting winds for EVA #3. EVA Team completed a post-EVA human factors survey on using the Mavic Pro standard drone controller.

Crew 207 identified the following useful items that could be 3D printed out of biodegradable, washable plastic at MDRS:

Headset ear piece

Look ahead plan : We will conduct a walking EVA tomorrow to continue our drone flight testing. We will continue to 3D print items at MDRS to save time and money for future crews. We will continue data collection for our thoracentesis and bike-powered 3D printing projects.

Anomalies in work : None

Weather : Sunny all day

Crew physical status : Fine

EVA :

Reports to be filed : Operations, GreenHab, Journalist, EVA Report, EVA Request

Support Requested : David to provide Crew 207 with a map showing the property boundaries of MDRS